OpenStudy (anonymous):

find vertex, line of symmetery, the max and min and graph. f(x)=-2x^2+2x+10

OpenStudy (anonymous):

-1/2(y-21/2)=(x-1/2)^2 ===> vertex=(1/2,21/2) , line of symmetry is x=1/2 , max point is (1/2,21/2)

OpenStudy (anonymous):